### v1

@@ 3.2 Ensemble approximation of transport filter
(16) $\boldsymbol{X}_{k}=\mathcal{T}_{k}(\hat{\boldsymbol{X}}_{k})=\hat{\boldsymbol{X}}_{k}+\mathbf{T}(\boldsymbol{Y}_{k}-\mathcal{H}(\hat{\boldsymbol{X}}_{k})),$
where $\mathbf{T}(\cdot)$ can be a linear or nonlinear map. Transport map $\mathcal{T}_{k}$ generates a transition probability that convert $\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k-1}}$ to $q^{\mathcal{T}_{k}}_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}$. Transport map $\mathcal{T}_{k}$ is related to the observation $\boldsymbol{y}_{k}$. We select a parameterized function space $\mathcal{G}_{0}$ and take map $\mathbf{T}$ to be the function in $\mathcal{G}_{0}$. Then $\mathcal{G}_{0}$ is transformed into space $\mathcal{G}$ according to Eq. 16. We now aims to find an optimal transport map $\mathcal{T}_{k}^{*}\in\mathcal{G}$ that minimizes the $\text{MMD}^{2}$ loss
(17) $\mathcal{T}_{k}^{*}=\underset{\mathcal{T}_{k}\in\mathcal{G}}{\arg\min}\operatorname{MMD}^{2}\left[\pi^{\mathcal{T}_{k}}_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}},\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}};\mathcal{F}\right],$
@@
In the framework of ensemble-based filter method, we start from an ensemble of particles $\{\boldsymbol{x}_{k-1}^{(i)}\}_{i=1}^{N}$ from the posterior density at time step $k-1$. Samples $\{\hat{\boldsymbol{x}}_{k}^{(i)}\}_{i=1}^{N}$ of the prior density at time step $k$ can be generated by the discrete dynamical model Eq. 1. Then the samples of prior density can be pushed forwards $q^{\mathcal{T}_{k}}_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}$ by transport map $\mathcal{T}_{k}$ in Eq. 16. In particle filtering, information on the true posterior density is implied in the weights of particles. We can express the reference posterior density as
(18) $\pi_{\boldsymbol{X}_{k}\mid\boldsymbol{Y}_{1:k}}\approx\sum_{i=1}^{N}w_{i}\delta_{\hat{\boldsymbol{x}}_{k}^{(i)}}(\boldsymbol{x}_{k}),w_{i}=\frac{\pi(\boldsymbol{y}_{k}\mid\hat{\boldsymbol{x}}_{k}^{(i)})}{\sum_{j=1}^{N}\pi(\boldsymbol{y}_{k}\mid\hat{\boldsymbol{x}}_{k}^{(j)})}.$
@@
(19) $\operatorname{MMD}^{2}=\sum_{i,j=1}^{m}$ $w_{i}w_{j}k(\hat{\boldsymbol{x}}_{k}^{(i)},\hat{\boldsymbol{x}}_{k}^{(j)})-2\sum_{i,j=1}^{m}w_{i}v_{j}k(\hat{\boldsymbol{x}}_{k}^{(i)},\mathcal{T}_{k}(\hat{\boldsymbol{x}}_{k}^{(j)}))$
$+\sum_{i,j=1}^{n}v_{i}v_{j}k(\mathcal{T}_{k}(\hat{\boldsymbol{x}}_{k}^{(i)}),\mathcal{T}_{k}(\hat{\boldsymbol{x}}_{k}^{(j)})),$
@@
[…]
@@
1:Given ensemble $\{\boldsymbol{x}_{k-1}^{(i)}\}_{i=1}^{N}$ at previous time step $k-1$, current observation $\boldsymbol{y}^{o}_{k}$, discrete dynamical model $\mathcal{M}_{k}(\cdot)$, observation operator $\mathcal{H}_{k}(\cdot)$, model uncertainty $\pi(\boldsymbol{\eta}_{k})$ and likelihood function $\pi(\boldsymbol{y}_{k}\mid\boldsymbol{x}_{k})$;
2:for $i=1,\cdots,N$ do
3:      $\hat{\boldsymbol{x}}_{k}^{(i)}=\mathcal{M}_{k}(\boldsymbol{x}_{k-1}^{(i)})+\boldsymbol{\eta}_{k}^{(i)}$;
4:end $\triangleright$ Forecast Step
5:Calculate particle weights $w_{i}=\frac{\pi(\boldsymbol{y}^{o}_{k}\mid\hat{\boldsymbol{x}}_{k}^{(i)})}{\sum_{j=1}^{N}\pi(\boldsymbol{y}^{o}_{k}\mid\hat{\boldsymbol{x}}_{k}^{(j)})}$;
6:Initialize transport map $\mathcal{T}_{k}$ defined as Eq. 16;
7:repeat
8:      for $i=1,\cdots,N$ do
9:           ${\boldsymbol{x}}_{k}^{(i)}=\mathcal{T}_{k}(\hat{\boldsymbol{x}}_{k}^{(i)})$; $\triangleright$ Alternative Analysis Step
10:           $v_{i}=1/N$;
11:      end for
12:      Calculate $\text{MMD}^{2}$ loss by Eq. 19;
13:      Update transport map $\mathcal{T}_{k}$ using gradient descent.
14:until Stopping criterion met
15:Approximated posterior ensemble $\{{\boldsymbol{x}}_{k}^{(i)}\}_{i=1}^{N}$;
3.3 Variance penalty for MMD loss
Given the nonlinear nature of the optimization problem, we may learn a local optimal transport map and vanilla MMD loss is less efficient in high dimensional problems. To make the transport map better approximate the posterior distribution, we introduce a variance penalty term. It prioritizes minimizing the discrepancy between the expectations of highly informative statistics for the approximated and reference posteriors. Since $\mathcal{F}$ is a unit ball of RKHS, the symmetry implies that
$\left[\underset{f\in\mathcal{F}}{\sup}\left(\mathbb{E}_{\boldsymbol{x}\sim p}[f(\boldsymbol{x})]-\mathbb{E}_{\boldsymbol{y}\sim q}[f(\boldsymbol{y})]\right)\right]^{2}=\underset{f\in\mathcal{F}}{\sup}\left(\mathbb{E}_{\boldsymbol{x}\sim p}[f(\boldsymbol{x})]-\mathbb{E}_{\boldsymbol{y}\sim q}[f(\boldsymbol{y})]\right)^{2}.$
@@ 3.3 Variance penalty for MMD loss
(20) $\text{Loss}=\underset{f\in\mathcal{F}}{\sup}\left\{\left(\mathbb{E}_{\boldsymbol{x}\sim p}[f(\boldsymbol{x})]-\mathbb{E}_{\boldsymbol{y}\sim q}[f(\boldsymbol{y})]\right)^{2}+\lambda\left(\mathbb{V}ar_{\boldsymbol{x}\sim p}[f(\boldsymbol{x})]+\mathbb{V}ar_{\boldsymbol{y}\sim q}[f(\boldsymbol{y})]\right)\right\},$
where $0\leq\lambda\leq 1$ is a regularization parameter. By a direct calculation, we can rewrite Eq. 20 as
@@
(21) $\text{Loss}=\underset{f\in\mathcal{F}}{\sup}\left\{(1-\lambda)\left(\mathbb{E}_{\boldsymbol{x}\sim p}[f(\boldsymbol{x})]-\mathbb{E}_{\boldsymbol{y}\sim q}[f(\boldsymbol{y})]\right)^{2}+\lambda\mathbb{E}_{\boldsymbol{x}\sim p,\atop\boldsymbol{y}\sim q}\left[f(\boldsymbol{x})-f(\boldsymbol{y})\right]^{2}\right\}.$
@@
In Eq. 21, training loss is split into two terms. The first term $\left(\mathbb{E}_{\boldsymbol{x}\sim p}[f(\boldsymbol{x})]-\mathbb{E}_{\boldsymbol{y}\sim q}[f(\boldsymbol{y})]\right)^{2}$ constrain the mean of two densities to be close, while the second term $\mathbb{E}\left[f(\boldsymbol{x})-f(\boldsymbol{y})\right]^{2}$ forces the distance of two related samples to be close. Two component terms of training loss are positive, and an upper bound estimate of Eq. 20 can be given by optimizing two component terms separately. The first term remains to be MMD, which can be interpreted as Eq. 15. Since $f\in\mathcal{F}$ and $\|f\|_{\mathcal{H}}\leq 1$, then
$\underset{f\in\mathcal{F}}{\sup}\mathbb{E}\left[f(\boldsymbol{x})-f(\boldsymbol{y})\right]^{2}$ $=\underset{f\in\mathcal{F}}{\sup}\mathbb{E}\left[\langle\phi_{\boldsymbol{x}}-\phi_{\boldsymbol{y}},f\rangle_{\mathcal{H}}^{2}\right]$

### v2

@@ 3.2 Ensemble approximation of transport filter
$\bm{X}_{k}=\mathcal{T}_{k}(\hat{\bm{X}}_{k})=\hat{\bm{X}}_{k}+\mathbf{T}(\bm{Y}_{k}-\mathbf{H}(\hat{\bm{X}}_{k})),$ (3.13)
where $\mathbf{T}(\cdot)$ can be a linear or nonlinear map. Transport map $\mathcal{T}_{k}$ generates a transition probability that convert $\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k-1}}$ to $q^{\mathcal{T}_{k}}_{\bm{X}_{k}\mid\bm{Y}_{1:k}}$ and is related to the observation $\bm{y}_{k}$. We select a parameterized function space $\mathcal{G}_{0}$ and take map $\mathbf{T}$ to be the function in $\mathcal{G}_{0}$. Then $\mathcal{G}_{0}$ is transformed into space $\mathcal{G}$ according to (3.13). We now aims to find an optimal transport map $\mathcal{T}_{k}^{*}\in\mathcal{G}$ that minimizes the $\text{MMD}^{2}$ loss
$\mathcal{T}_{k}^{*}=\underset{\mathcal{T}_{k}\in\mathcal{G}}{\arg\min}\operatorname{MMD}^{2}\left[\pi^{\mathcal{T}_{k}}_{\bm{X}_{k}\mid\bm{Y}_{1:k}},\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k}};\mathcal{F}\right],$ (3.14)
@@
In the framework of ensemble-based filter method, we start from an ensemble of particles $\{\bm{x}_{k-1}^{(i)}\}_{i=1}^{N}$ from the posterior density at time step k1k-1. Samples $\{\hat{\bm{x}}_{k}^{(i)}\}_{i=1}^{N}$ of the prior density at time step $k$ can be generated by the discrete dynamical model (2.1). Then the samples of prior density can be pushed forwards $q^{\mathcal{T}_{k}}_{\bm{X}_{k}\mid\bm{Y}_{1:k}}$ by transport map $\mathcal{T}_{k}$ in (3.13). In particle filtering, information on the true posterior density is implied in the weights of particles. We can express the reference posterior density as
$\pi_{\bm{X}_{k}\mid\bm{Y}_{1:k}}\approx\sum_{i=1}^{N}w_{i}\delta_{\hat{\bm{x}}_{k}^{(i)}}(\bm{x}_{k}),w_{i}=\frac{\pi(\bm{y}_{k}\mid\hat{\bm{x}}_{k}^{(i)})}{\sum_{j=1}^{N}\pi(\bm{y}_{k}\mid\hat{\bm{x}}_{k}^{(j)})}.$ (3.15)
@@
$\operatorname{MMD}^{2}=\sum_{i,j=1}^{m}$ $w_{i}w_{j}k(\hat{\bm{x}}_{k}^{(i)},\hat{\bm{x}}_{k}^{(j)})-2\sum_{i,j=1}^{m}w_{i}v_{j}k(\hat{\bm{x}}_{k}^{(i)},\mathcal{T}_{k}(\hat{\bm{x}}_{k}^{(j)}))$ (3.16)
$+\sum_{i,j=1}^{n}v_{i}v_{j}k(\mathcal{T}_{k}(\hat{\bm{x}}_{k}^{(i)}),\mathcal{T}_{k}(\hat{\bm{x}}_{k}^{(j)})),$
@@
The selection of the kernel function in the ensemble transport filter is crucial in determining how well the ensemble transport filter approximates the posterior distribution. When we choose a universal kernel such as the Gaussian kernel $k(\bm{x},\bm{x}^{\prime})=\exp\left(-\|\bm{x}-\bm{x}^{\prime}\|/r^{2}\right)$, ensemble transport filter will tend to approximate the posterior distribution as accurate as possible. When we choose a linear kernel function $k(\bm{x},\bm{x}^{\prime})=\bm{x}^{\top}\bm{x}^{\prime}+1$, it tends to obtain a better estimate of the mean of the posterior distribution and ignores higher order moment information. For further analysis, one can adaptively adjust the kernel functions based on the actual model, as discussed in [33].
@@
0: Given ensemble $\{\bm{x}_{k-1}^{(i)}\}_{i=1}^{N}$ at previous time step k1k-1, current observation $\bm{y}^{o}_{k}$, discrete dynamical model $\mathbf{M}_{k}(\cdot)$, observation operator $\mathbf{H}_{k}(\cdot)$, model uncertainty $\pi(\bm{\eta}_{k})$ and likelihood function $\pi(\bm{y}_{k}\mid\bm{x}_{k})$;
1:for $i=1,\cdots,N$ do
2:  $\hat{\bm{x}}_{k}^{(i)}=\mathbf{M}_{k}(\bm{x}_{k-1}^{(i)})+\bm{\eta}_{k}^{(i)}$;
3:end for{Forecast Step}
4: Calculate particle weights $w_{i}=\frac{\pi(\bm{y}^{o}_{k}\mid\hat{\bm{x}}_{k}^{(i)})}{\sum_{j=1}^{N}\pi(\bm{y}^{o}_{k}\mid\hat{\bm{x}}_{k}^{(j)})}$;
5: Initialize transport map $\mathcal{T}_{k}$ defined as (3.13);
6:repeat
7:  for $i=1,\cdots,N$ do
8:   ${\bm{x}}_{k}^{(i)}=\mathcal{T}_{k}(\hat{\bm{x}}_{k}^{(i)})$; {Alternative Analysis Step}
9:   $v_{i}=1/N$;
10:  end for
11:  Calculate $\text{MMD}^{2}$ loss by (3.16);
12:  Update transport map $\mathcal{T}_{k}$ using gradient descent.
13:until Stopping criterion met
13: Approximated posterior ensemble $\{{\bm{x}}_{k}^{(i)}\}_{i=1}^{N}$;
3.2.1 Variance penalty for MMD loss
Given the nonlinear nature of the optimization problem, the learned transport map may be locally optimal. Additionally, vanilla MMD loss is less efficient in high dimensional problems. To make the transport map better approximate the posterior distribution, we introduce a variance penalty term. It prioritizes minimizing the discrepancy between the expectations of highly informative statistics for the reference posteriors. Suppose $p$ is the reference distribution and $q$ the approximation posterior. Since $\mathcal{F}$ is a unit ball of RKHS, the symmetry implies that
$\left[\underset{f\in\mathcal{F}}{\sup}\left(\mathbb{E}_{\bm{x}\sim p}[f(\bm{x})]-\mathbb{E}_{\bm{y}\sim q}[f(\bm{y})]\right)\right]^{2}=\underset{f\in\mathcal{F}}{\sup}\left(\mathbb{E}_{\bm{x}\sim p}[f(\bm{x})]-\mathbb{E}_{\bm{y}\sim q}[f(\bm{y})]\right)^{2}.$
@@ 3.2.1 Variance penalty for MMD loss
$\text{Loss}=\underset{f\in\mathcal{F}}{\sup}\left\{\left(\mathbb{E}_{\bm{x}\sim p}[f(\bm{x})]-\mathbb{E}_{\bm{y}\sim q}[f(\bm{y})]\right)^{2}+2\mathbb{V}ar_{\bm{x}\sim p}[f(\bm{x})]\right\}.$ (3.17)
The penalty term does not depend on the approximate distribution. As a result, when minimizing the loss function, it does not compromise the ability of the MMD term to enforce accurate distributional matching between the approximated distribution $q$ and the reference $p$. By a direct calculation, we can rewrite (3.17) as
@@
$\operatorname{Loss}$ $=\underset{f\in\mathcal{F}}{\sup}\left\{\left(\mathbb{E}_{\bm{x}\sim p,\atop\bm{y}\sim q}\left[f(\bm{x})-f(\bm{y})\right]^{2}\right)+\left(\mathbb{V}ar_{\bm{x}\sim p}[f(\bm{x})]-\mathbb{V}ar_{\bm{y}\sim q}[f(\bm{y})]\right)\right\}$ (3.18)
$\leq\underset{f\in\mathcal{F}}{\sup}\ \mathbb{E}_{\bm{x}\sim p,\atop\bm{y}\sim q}\left[f(\bm{x})-f(\bm{y})\right]^{2}+\underset{f\in\mathcal{F}}{\sup}\ \left(\mathbb{V}ar_{\bm{x}\sim p}[f(\bm{x})]-\mathbb{V}ar_{\bm{y}\sim q}[f(\bm{y})]\right),$
@@
the equality holds when $p=q$. In (3.18), training loss is split into two terms. The first term $\mathbb{E}_{\bm{x}\sim p,\atop\bm{y}\sim q}\left[f(\bm{x})-f(\bm{y})\right]^{2}$ constrain the expectation of the $L^{2}$-distance between the samples drawn from two densities to be close, while the second term $\left(\mathbb{V}ar_{\bm{x}\sim p}[f(\bm{x})]-\mathbb{V}ar_{\bm{y}\sim q}[f(\bm{y})]\right)$ forces the discrepancy of the variance of two distribution to be close. Since $f\in\mathcal{F}$ and $\|f\|_{\mathcal{H}}\leq 1$, then
$\underset{f\in\mathcal{F}}{\sup}\mathbb{E}\left[f(\bm{x})-f(\bm{y})\right]^{2}$ $=\underset{f\in\mathcal{F}}{\sup}\mathbb{E}\left[\langle\phi_{\bm{x}}-\phi_{\bm{y}},f\rangle_{\mathcal{H}}^{2}\right]$